OpenStudy (anonymous):

The king and Parliament viewed the American colonies as a a. place for vacations. c. source of irritation. b. drain on their economy. d. source of funds.

OpenStudy (jack1):

\[\Large -\frac {\sqrt{(a^2-x^2)}}{x^2}~~+~~\frac {1}{a \sqrt{1-\frac {x^2}{a^2}}}~~-~~ \frac {1}{\sqrt{a^2-x^2}}\] \[\Large -\sqrt{(a^2-x^2)}~~+~~\frac {x^2}{a \sqrt{1-\frac {x^2}{a^2}}}~~-~~ \frac {x^2}{\sqrt{a^2-x^2}}\] \[\Large -\sqrt{a^2-x^2}~~\times~~\sqrt{a^2-x^2}~~+~~\frac {x^2~~\times~~\sqrt{a^2-x^2}}{a \sqrt{1-\frac {x^2}{a^2}}}~~-~~ x^2\] \[\large -\sqrt{a^2-x^2}~~\times~~\sqrt{a^2-x^2}~~ \times~~ a~~+~~\frac {x^2~~\times~~\sqrt{a^2-x^2}}{ \sqrt{1-\frac {x^2}{a^2}}}~~-~~ x^2 \times a\] \[\large -(a^2-x^2)~~ \times~~ a~~+~~\frac {x^2~~\times~~\sqrt{a^2-x^2}}{ \sqrt{1-\frac {x^2}{a^2}}}~~-~~ x^2 \times a\] \[\large -a^3+ax^2~~+~~\frac {x^2~~\times~~\sqrt{a^2-x^2}}{ \sqrt{1-\frac {x^2}{a^2}}}~~-~~ ax^2\] \[\large -a^3+ax^2 - ax^2~~+~~\frac {x^2~~\times~~\sqrt{a^2-x^2}}{ \sqrt{1-\frac {x^2}{a^2}}}\] \[\large -a^3~~+~~\frac {x^2~~\times~~\sqrt{a^2-x^2}}{ \sqrt{1-\frac {x^2}{a^2}}}\] \[\large (-a^3~~\times~~{ \sqrt{1-\frac {x^2}{a^2}})~~+~~ ({x^2~~\times~~\sqrt{a^2-x^2}}})\]